OpenStudy (anonymous):

solve each of the quadratic equations below and describe what the solutions represent to the graph of each: y=x^2 + 3x + 2 y=x^2 +2x + 1

OpenStudy (anonymous):

help

OpenStudy (anonymous):

Factor each function and check for which x values it equals 0.

OpenStudy (anonymous):

for the first equation i am getting x= -1/4 Can that be correct?

OpenStudy (mathstudent55):

Are these two separate problems or is this a system of quadratic equations?

OpenStudy (anonymous):

Separate problems @mathstudent55

OpenStudy (anonymous):

two separate problems to describe what solutions represent to the graph of each. I got X= -1/4 and I am not sure that can be correct.

OpenStudy (anonymous):

\[y=x^2+3x+2 \rightarrow y=(x+1)(x+2) \rightarrow (x+1)(x+2)=0 \implies x = -1,-2\]Get what I did here? Now try doing the second one. @mathgirl2012

OpenStudy (mathstudent55):

Ok, set each right side equal to zero. y=x^2 + 3x + 2 So do x^2 + 3x + 2 = 0 Factor the left side, and solve for x.

OpenStudy (anonymous):

Having any success? @mathgirl2012

OpenStudy (anonymous):

I did that and got x=1 x=2 then I solved for the vertex x=-b/2a and I got x+-1/4?

OpenStudy (mathstudent55):

Then do the same for the second equation: y=x^2 +2x + 1 Set x^2 + 2x + 1 = 0 Factor the left side, and solve for x.

OpenStudy (mathstudent55):

For the first one, the solutions are x = -2 or x = -1

OpenStudy (anonymous):

second equation i got x=1

OpenStudy (mathstudent55):

For first eq, -b/2a = -3/2 = -1.5

OpenStudy (mathstudent55):

The vertex is at x = -1.5

OpenStudy (mathstudent55):

After factoring, be careful. The firsrt equation factors into (x + 2)(x + 1) = 0 Next step is: x + 2 = 0 or x + 1 = 0, so you get x = -2 or x = -1 The answers are NEGATIVE 2 and NEGATIVE 1, not 2 and 1

OpenStudy (anonymous):

Second equation is:\[y=x^2+2x+1 \rightarrow y = (x+1)^2 \rightarrow (x+1)^2=0 \implies x = -1\]

OpenStudy (mathstudent55):

Read my explanation just above.

OpenStudy (anonymous):

For the first equation I got x=1, x=2 how did you get x=-2 , x= -1 when the equation reads x^2 + 3x + 2??? @mathstudent55

OpenStudy (anonymous):

Thank you

OpenStudy (mathstudent55):

@mathgirl2012 I just answered. Read my post just a little above.

OpenStudy (anonymous):

Because when you factor it, you get y = (x + 1)(x + 2). Then when you are finding solutions, you set the it to equal 0 like this: (x + 1)(x + 2) = 0. Now you find x-values where the function is 0. You know that if the first bracket is 0, then the whole thing is 0, if the second bracket 0, then the whole function is also. So if x + 1 = 0, then x = -1. If x + 2 = 0, then x = -2. Therefore, the zeroes are x = -1, -2 because at these values, the whole function becomes 0. @mathgirl2012

OpenStudy (mathstudent55):

x^2 + 3x + 2 = 0 (x + 2)(x + 1) = 0 x + 2 = 0 or x + 1 = 0 x = -2 or x = -1

OpenStudy (mathstudent55):

Now for the vertex, which has x = -b/2a, use the a, b, and c of the original equation. You have x^2 + 3x + 2 = 0 When comparred with ax^2 + bx + c = 0, you see that a = 1, b = 3, and c = 2, so -a/2b = -2/3 = -1.5 The vertex has an x-coordinated of -1.5

OpenStudy (anonymous):

your the best you helped me and i wish i could give you both medals @mathstudent55 and @genius12 

OpenStudy (anonymous):

Awwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwwww. You can Fan me =D lol

OpenStudy (mathstudent55):

Thanks. I'll take care of genius12

OpenStudy (anonymous):

lol <3 @mathstudent55

OpenStudy (anonymous):

@genius12 and @mathstudent55 you guys can fan me 2 lol ;P